OpenStudy (anonymous):

When a toy truck collides into a toy car, the momentum of ____ is the same before and after the collision. a. the truck c. the truck times the car b. the car d. the truck plus the car

OpenStudy (amistre64):

is it an elastic or inelastic collision?

OpenStudy (amistre64):

what were the speeds to begin with?

OpenStudy (anonymous):

Idk Thats Just The Question

OpenStudy (amistre64):

well there are formulas for these things

OpenStudy (amistre64):

mv1 + mv2 = mV1 + mV2 for elastic mv1 + mv2 =v3(m1+m2) for inelastic

OpenStudy (amistre64):

truck plus car is what i think they are going after

OpenStudy (amistre64):

the sum of the parts is the same